OpenStudy (anonymous):

Describe, in your own words, the identity sin^2(x) + cos^2(x) = 1.

OpenStudy (goformit100):

sin^2(x) + cos^2(x) = 1

OpenStudy (goformit100):

we know:- sin(x) = perpendicular/hypotenuse = p/h cos(x)= base/hypotenuse = b/h Now,